# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

### v6

@@ Abstract
Future Frame Synthesis (FFS) focuses on generating future frame sequences conditioned on existing content. This survey provides a comprehensive review of existing research on FFS, covering commonly used datasets and representative algorithms. We discuss key challenges and trace the evolution of FFS in computer vision, particularly the shift from deterministic to generative approaches. Our taxonomy outlines major advances and methodological shifts, emphasizing the rising significance of generative models in producing realistic and diverse predictions. This survey offers a comprehensive analysis of current research and, moreover, suggests promising avenues for future exploration in this ever-changing domain.
@@ 1 Introduction
As a foundation, we introduce the problem formulation and core challenges in Section 2. Our taxonomy is organized around the degree of stochasticity in modeling approaches. The field of FFS is rapidly moving from traditional deterministic prediction to large-scale generative approaches as the latest video generation models continue to emerge. These models have made important advances in parameter scales, generation lengths, control capabilities, and training strategies, injecting new vitality. In Table 2, we present deterministic algorithms that aim to perform pixel-level fitting based on fixed target frames. However, pixel-level metrics tend to drive models to average over multiple plausible futures, often resulting in blurry predictions. In Section 4, we examine algorithms that enable stochastic motion prediction. These include approaches that inject stochastic variables into deterministic models, as well as methods based on explicit probabilistic modeling. Such methods enable sampling from motion distributions, yielding diverse yet plausible predictions beyond the target frame. Given current FFS algorithms’ limited generative capacity, especially for high-resolution videos involving object appearance and disappearance, we discuss the generative FFS methods in Section 5. These methods prioritize producing coherent long-term sequences over pixel-level accuracy. In Section 6, we explore the broad applicability of FFS in areas such as world model, autonomous driving, robotics, film production, meteorology, and anomaly detection. These use cases demonstrate the role of FFS in dynamic scene understanding and interaction. In Section 7, we review prior surveys on video prediction and diffusion-based video generation. We also clarify our distinct focus: a comprehensive analysis of FFS spanning from deterministic to generative paradigms, emphasizing the growing role of generative models in producing realistic and diverse future predictions.
@@ 2.2 Paradigms and Architectures
The deterministic paradigm emphasizes pixel-level fitting to fixed target frames, typically employing low-level computer vision architectures such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012), Recurrent Neural Networks (RNNs) (Rumelhart et al., 1985), Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and U-Net (Ronneberger et al., 2015). Recently, the Transformer architecture has gradually challenged the dominance of traditional CNNs and RNNs. Vision Transformer (ViT) (Dosovitskiy et al., 2021) divide images into fixed-size patches and treat these patches as sequences input into the Transformer for processing. Swin Transformer (Liu et al., 2021) introduces a hierarchical Transformer structure that handles features at different scales through local window self-attention mechanisms and progressively expanding receptive fields. In low-level vision tasks, IPT (Chen et al., 2021) employs pre-trained Transformers to address various low-level visual tasks. Models like TimeSformer (Bertasius et al., 2021) and ViViT (Arnab et al., 2021) utilize Transformers to process spatio-temporal information in videos by combining video frames and time steps.
@@ 2.3.1 Learning objectives and evaluation metrics
Regarding the quality assessment of video generation models, some more comprehensive evaluation systems have been recently established. VBench (Huang et al., 2024) benchmark suite has been proposed to address these concerns by providing a comprehensive evaluation framework for video generative models. VBench evaluates video generative models comprehensively, covering perceptual quality, dynamics, temporal consistency, content diversity, and prompt alignment for a holistic performance assessment. FVMD (Liu et al., 2024a) designs explicit motion features based on key point tracking, focusing on evaluating motion consistency in video generation. VBench-2.0 (Zheng et al., 2025) further focuses on the intrinsic faithfulness of generative models, such as whether they adhere to real-world principles, including physical laws and commonsense reasoning. As people’s pursuit of generating realism gradually increases, the evaluation of real physical interaction and real-scene simulation will be a promising research topic. Moreover, for FFS, we need to consider whether the generated results are reasonable given the observed results.
@@ 2.3.3 Generalization
Figure 1: Datasets. Datasets from KTH Action (Schuldt et al., 2004), MovingMNIST (Srivastava et al., 2015), Something-Something (Goyal et al., 2017), Vimeo-90K (Xue et al., 2019), WebVid-10M (Bain et al., 2021), DropletVideo-10M (Zhang et al., 2025), BDD100K (Yu et al., 2020). We adjust the absolute resolution of the image for visualization while maintaining the relative size relationships.
@@
ChronoMagic (Yuan et al., 2025) General 2, 265 variable $1920\times 1080$ Text
@@
Pandas-70M (Chen et al., 2024b) General 70, 000, 000 variable $1280\times 720$ Class, Text
@@
ConsisID (Yuan et al., 2024) Human 54, 239 129 $1280\times 720$ Text
@@ 2.4 Datasets
The advancement of video synthesis models heavily depends on the diversity, quality, and characteristics of training datasets. A common observation is that the suitability of datasets varies with their dimensionality and size: lower-dimensional datasets, often smaller in scale, tend to exhibit limited generalization. In contrast, higher-dimensional datasets offer greater variability, contributing to stronger generalization capabilities. In general, video complexity and resolution increase in recent years, as shown in Figure 1. In Table 1, we summarize the most widely used datasets in video synthesis, highlighting their scale and available supervisory signals to provide a comprehensive overview of the current dataset landscape. For datasets lacking detailed reports in their original papers or project pages, we estimate missing values using mean or median statistics to ensure consistency across the analysis.
@@ 3.1 Raw Pixel Space
In short-term FFS, approaches operating in the raw pixel space have achieved promising results. In this section, we review representative methods and discuss the associated challenges.
@@ 3.1.3 Motion-based synthesis
There are still ample opportunities for exploration in the representation of motion. For example, GaussianPrediction Zhao et al. (2024) explores the use of 3D Gaussian representations for modeling the appearance and geometry of dynamic scenes, enabling the rendering of images for future scenarios. MoSca (Lei et al., 2024) proposes the Motion Scaffold representation, which lifts the predictions of 2D foundational models (such as depth estimation, pixel trajectories, and optical flow) to 3D and optimizes them using physics-inspired constraints. Ultimately, MoSca reconstructs the geometry and appearance of the scene with a set of dynamic Gaussians (Wu et al., 2024b). Seamlessly integrating new representation methods with existing frameworks presents novel challenges.
3.1.4 Transformers
Challenges.
@@ 4 Stochastic Synthesis
In this section, we explore stochastic synthesis methods, including GANs and VAEs, which were originally categorized for their emphasis on modeling randomness and uncertainty in video motion. Although these models possess generative capabilities, they were originally more closely associated with stochasticity due to their primary focus on capturing the inherent variability and unpredictability of video sequences. Current research prefers to simulate the complex dynamics of the future using end-to-end generative modeling strategies rather than by introducing stochastic distributions or probabilistic models into deterministic algorithms to obtain diverse outcomes. In Section 5, we discuss generative synthesis approaches, such as diffusion models and auto-regressive models, which are explicitly designed to prioritize the generation of diverse, high-quality video content. This distinction reflects the evolving focus of video synthesis research, shifting from an emphasis on stochasticity to a broader emphasis on generative capability.
@@ 4.3 Motion-Controllable Synthesis
SEINE (Chen et al., 2024d) ICLR’24 Propose a short-to-long video diffusion model, aimed at generating coherent long videos at the "story-level".
@@
CineMaster (Wang et al., 2025a) preprint A framework for 3D-aware and controllable text-to-video generation, empowering users with precise control over the scene.
@@
LivePhoto (Chen et al., 2024c) ECCV’24 Enable users to control the temporal motion of images with text, and decode motion-related instructions into videos using a generator equipped with a motion module.
@@
LVT (Rakhimov et al., 2021) VISIGRAPP’21 Reduce the computational requirements of video generation models by modeling dynamics in the latent space.
@@ 5 Generative Synthesis
The emergence and disappearance of objects introduce unpredictability, and generative models require a profound understanding of the underlying physical principles that govern the real world. Rather than relying on simplistic linear motion predictions extrapolated from historical frames, they address the challenge using sophisticated and imaginative modeling techniques. As a result, tasks such as transforming one static image into one dynamic video—often referred to as the image animation problem—have emerged as promising candidates for applying generative video prediction techniques.
@@ 5.1 Diffusion-Based Generation
Extensions of LDM have demonstrated strong generative capabilities in video systhesis (Voleti et al., 2022). For instance, Video LDM (Blattmann et al., 2023b) leverages pre-trained image models to generate videos, enabling multi-modal, high-resolution, and long-term video synthesis. Similarly, SEINE (Chen et al., 2024d) introduces a versatile video diffusion model capable of generating transition sequences, thereby extending short clips into longer videos. Burgert et al. (2025) presents a scalable method for finetuning video diffusion models to control motion without modifying model architectures or training pipelines. It introduces a noise warping algorithm that replaces random temporal Gaussian noise with optical flow-guided warped noise, while maintaining spatial Gaussianity. Magic141 (Yi et al., 2025) decomposes the text-to-video generation task into two simpler subtasks for diffusion step distillation: text-to-image generation and image-to-video generation. It shows that image-to-video generation converges more easily than direct text-to-video generation under the same optimization setup, and explores the trade-off between computational cost and video quality.
Text-guided video completion with additional information.
@@
Recent research efforts have focused on harnessing additional modalities alongside RGB images to accomplish the task of text-guided video completion. LFDM (Ni et al., 2023) extends latent diffusion models to synthesize optical flow sequences in the latent space based on textual guidance. Seer (Gu et al., 2024) inflates Stable Diffusion (Rombach et al., 2022) along the temporal axis, enabling the model to utilize natural language instructions and reference frames to envision multiple variations of future outcomes. Emu Video (Girdhar et al., 2024) generates an image conditioned on textual guidance and extrapolates it into a video, making it adaptable to diverse textual inputs. DynamiCrafter (Xing et al., 2024) extends text-guided image animation to open-domain image scenarios. SparseCtrl (Guo et al., 2024) supports sketch-to-video generation, depth-to-video generation, and video prediction with an expanded range of input modalities. Other methods, such as PEEKABOO (Jain et al., 2024), explore interactive synthesis, aiming to unlock unprecedented applications and creative potential. VACE (Jiang et al., 2025) proposes a unified conditioning format called the video condition unit, which encodes text instructions, reference images, and masks. It incorporates a context adapter module that injects task-specific context into the model using structured temporal and spatial representations. This design ensures coherence across frames and spatial regions during generation and supports diverse applications through flexible task composition. SkyReels-A2 (Fei et al., 2025) defines the elements-to-video task, which involves generating videos by integrating arbitrary visual elements such as characters, objects, and backgrounds through reference images under the guidance of text prompts. The model demonstrates a strong ability to preserve the distinct visual details of each individual element throughout the generation process while simultaneously producing results that are coherently blended. This capability represents a significant step forward in the field by shifting the focus of video synthesis from single-condition generation to multi-condition combinatorial control. CineMaster (Wang et al., 2025a) proposes a 3D-aware, controllable text-to-video framework where users interactively define 3D control signals such as object bounding boxes and camera trajectories to guide a text-to-video diffusion model. To address the scarcity of annotated data, it introduces an automated pipeline that extracts 3D control signals from large-scale video datasets.
@@
Some works aim to achieve a more precise interpretation of textual guidance and preserve this information across the temporal dimension. MicroCinema (Wang et al., 2024b) adopts a divide-and-conquer strategy to address challenges related to appearance and temporal coherence. It employs a two-stage generation pipeline, first creating an initial image using an existing text-to-image generator, and then introducing a dedicated text-guided video generation framework for motion modeling. LivePhoto (Chen et al., 2024c) proposes a framework that incorporates motion intensity as an auxiliary factor to enhance control over motion dynamics. It further introduces a text re-weighting mechanism to emphasize motion-related descriptions, demonstrating strong performance in text-guided video synthesis tasks. I2VGen-XL (Zhang et al., 2023) utilizes static images to provide semantic and quality-related guidance, highlighting the diversity of approaches in text-guided video synthesis. SkyReels-V2 (Chen et al., 2025) introduces a multi-modal LLM to interpret narrative semantics by decomposing script-like prompts into shot-level descriptions and utilizes a specialized video captioning module to automatically annotate training data with rich cinematographic language. This methodology enables its 14B parameter model to generate videos exceeding 30 seconds in duration while exhibiting a professional cinematic style, representing the first open-source model to achieve long-form video synthesis that closely approximates the quality and coherence of film production.
@@
Despite the rising dominance of diffusion models in generative tasks, some researchers seek to preserve the architecture of auto-regressive models to simultaneously leverage the strengths of both paradigms (Weng et al., 2024). Early generative video synthesis algorithms are constrained by limited data availability and model scalability, yet they lay the groundwork for integrating LLMs and diffusion models in video generation. VDT (Lu et al., 2024) represents one of the earliest efforts to incorporate a Transformer-based backbone into diffusion models for video generation. Inspired by the success of DiT (Peebles & Xie, 2023) in image synthesis, VDT replaces the conventional U-Net with a spatio-temporal Transformer that explicitly models both spatial and temporal dependencies. W.A.L.T. (Gupta et al., 2025) provides the first successful empirical demonstration of a Transformer-based backbone for jointly training image and video latent diffusion models. Recently, GAIA-1 (Hu et al., 2023a) and Sora (Brooks et al., 2024) have also leveraged the strengths of DiT to enable more creative, generalizable, and scalable video synthesis. RIVER (Davtyan et al., 2023) leverages flow matching (Lipman et al., 2023) for efficient video prediction by conditioning on a small set of past frames in the latent space of a pre-trained VQGAN (Esser et al., 2021). Collectively, these advances underscore the potential of diffusion models to produce high-quality, controllable, and diverse video content, thereby pushing the boundaries of what is achievable in computer vision and artificial intelligence.
@@
HunyuanVideo (Kong et al., 2024) is a very large open-source video generation model (13B parameters). Relying on key strategies such as data curation and image-video co-training, the model reaches a level of generation quality comparable to or even better than that of the leading closed-source models, which effectively narrows the performance gap between open-source and closed-source solutions.  Ma et al. (2025) release an even larger video generation model, Step-Video-T2V with 30B parameters, and a new benchmark for video generation, Step-Video-T2V-Eval. The model incorporates a deep compression video Variational Autoencoder for compact and efficient video representation, and leverages a flow matching approach to train a DiT (Peebles & Xie, 2022) architecture for denoising input noises into latent frames. To enhance the visual quality and mitigate artifacts, a video-based direct preference optimization method is applied. For the task of text-driven image-to-video generation, Huang et al. (2025) release Step-Video-TI2V, a model of equal scale with 30B parameters, along with its corresponding benchmark, Step-Video-TI2V-Eval. It is noteworthy that improving video generation performance does not necessarily require increasing the number of model parameters. Seaweed-7B (Seawead et al., 2025), with only 7B parameters, demonstrates generation quality comparable to that of models with tens of billions of parameters. This provides important insights into how the performance of medium-sized DiT models can be effectively enhanced without relying solely on scaling up model size. Sand-AI (2025) has open-sourced a 24B parameter video generation model, MAGI-1, that divides video frame sequences into fixed-length video chunks and adopts a chunk-wise auto-regressive diffusion framework, facilitating text-driven image-to-video generation through chunk-wise prompting. This design illustrates a promising direction in which high-fidelity video synthesis and fine-grained instruction control can be unified.
@@
While diffusion models have made significant strides in video generation, several critical challenges remain. Ensuring temporal coherence and consistency across frames is essential for achieving realism, yet remains a major challenge (Chen et al., 2024e; Xu et al., 2024). The computational efficiency and scalability of diffusion models are hindered by their resource-intensive nature, which limits their broader adoption (Peebles & Xie, 2023). Issues of controllability and interpretability persist, as textual guidance does not always align with visual outcomes, and model behaviors often remain opaque. Data availability and diversity are critical for training robust models, but acquiring comprehensive and diverse datasets remains a major bottleneck.

### v7

@@ Abstract
Future Frame Synthesis (FFS) is dedicated to generating future frame sequences based on existing content, expanding the scope of Video Frame Prediction by emphasizing the synthesis aspect. This survey provides a comprehensive review of existing research on FFS, covering commonly used datasets and representative algorithms. We discuss key challenges and trace the evolution of FFS in computer vision, particularly the shift from deterministic to generative approaches. Our taxonomy outlines major advances and methodological shifts, emphasizing the rising significance of generative models in producing realistic and diverse predictions. This survey offers a comprehensive analysis of current research and, moreover, suggests promising avenues for future exploration in this ever-changing domain.
@@ 1 Introduction
The field of FFS is rapidly moving from traditional deterministic prediction to large-scale generative approaches as the latest video generative models continue to emerge. These models have made important advances in parameter scales, generation lengths, control capabilities, and training strategies, injecting new vitality. As a foundation, we introduce the problem formulation and core challenges in Section 2. Our taxonomy is organized around the degree of stochasticity in modeling approaches. In Table 2, we present deterministic algorithms that aim to perform pixel-level fitting based on fixed target frames. However, pixel-level metrics tend to drive models to average over multiple plausible futures, often resulting in blurry predictions. In Section 4, we examine algorithms that enable stochastic motion prediction. These include approaches that inject stochastic variables into deterministic models, as well as methods based on explicit probabilistic modeling. Such methods enable sampling from motion distributions, yielding diverse yet plausible predictions beyond the target frame. Given current FFS algorithms’ limited generative capacity, especially for high-resolution videos involving object appearance and disappearance, we discuss the generative FFS methods in Section 5. These methods prioritize producing coherent long-term sequences over pixel-level accuracy. In Section 6, we explore the broad applicability of FFS in areas such as world model, autonomous driving, robotics, film production, meteorology, and anomaly detection. These use cases demonstrate the role of FFS in dynamic scene understanding and interaction. In Section 7, we review prior surveys on video prediction and diffusion-based video generation. We also clarify our distinct focus: a comprehensive analysis of FFS spanning from deterministic to generative paradigms, emphasizing the growing role of generative models in producing realistic and diverse future predictions.
@@ 2.2 Paradigms and Architectures
The deterministic paradigm emphasizes pixel-level fitting to fixed target frames, typically employing low-level computer vision architectures such as Convolutional Neural Networks (CNNs) (Krizhevsky et al., 2012), Recurrent Neural Networks (RNNs) (Rumelhart et al., 1985), Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and U-Net (Ronneberger et al., 2015). Recently, the Transformer architecture has gradually challenged the dominance of traditional CNNs and RNNs. A landmark work, Vision Transformer (ViT) (Dosovitskiy et al., 2021), divides images into fixed-size patches and treats these patches as sequences input into the Transformer for processing. Swin Transformer (Liu et al., 2021) further introduces a hierarchical Transformer structure that handles features on different scales through a local window self-attention mechanism and progressively expanding receptive fields. In low-level vision tasks, IPT (Chen et al., 2021) employs pre-trained Transformers to address various low-level visual tasks. And Models like TimeSformer (Bertasius et al., 2021) and ViViT (Arnab et al., 2021) utilize Transformers to process spatio-temporal information in videos by combining video frames and time steps.
@@ 2.3.1 Learning objectives and evaluation metrics
Regarding the quality assessment of video generative models, some more comprehensive evaluation systems have been recently established. VBench (Huang et al., 2024) benchmark suite has been proposed to address these concerns by providing a comprehensive evaluation framework for video generative models. VBench evaluates video generative models comprehensively, covering perceptual quality, dynamics, temporal consistency, content diversity, and prompt alignment for a holistic performance assessment. FVMD (Liu et al., 2024a) designs explicit motion features based on key point tracking, focusing on evaluating motion consistency in video generation. VBench-2.0 (Zheng et al., 2025) further focuses on the intrinsic faithfulness of generative models, such as whether they adhere to real-world principles, including physical laws and common sense reasoning. Recently, some of the work has increasingly utilized multi-modal LLMs to evaluate the generated results. For example, ViStoryBench (Zhuang et al., 2025) leverages GPT-4o (Hurst et al., 2024) to assess the instruction-following ability generative models from multiple aspects. As people’s pursuit of generating realism gradually increases, the evaluation of real physical interaction and real-scene simulation will be a promising research topic. Moreover, for FFS, we need to consider whether the generated results are reasonable given the observed results.
@@ 2.3.3 Generalization
Figure 1: Dataset samples. We adjust the absolute resolution of the image for visualization while maintaining the relative size relationships.
@@
ChronoMagic (Yuan et al., 2025b) General 2, 265 variable $1920\times 1080$ Text
@@
Pandas-70M (Chen et al., 2024c) General 70, 000, 000 variable $1280\times 720$ Class, Text
@@
ConsisID (Yuan et al., 2025a) Human 54, 239 129 $1280\times 720$ Text
@@ 2.4 Datasets
The advancement of video synthesis models is highly dependent on the diversity, quality, and characteristics of training datasets. A common observation is that the suitability of datasets varies with their dimensionality and size: lower-dimensional datasets, often smaller in scale, tend to exhibit limited generalization. Conversely, higher-dimensional datasets offer greater variability, contributing to stronger generalization capabilities. Generally, video complexity and resolution have increased in recent years, as shown in Figure 1. In Table 1, we summarize the most widely used datasets in video synthesis, highlighting their scale and available supervisory signals to provide a comprehensive overview of the current dataset landscape. For datasets lacking detailed reports in their original papers or project pages, we estimate missing values using mean or median statistics to ensure consistency across the analysis.
@@ 3.1 Raw Pixel Space
In short-term FFS, approaches operating in the raw pixel space have achieved promising results. In this section, we review representative methods and discuss the associated challenges. We will first introduce the evolution of the architecture of popular backbones, and then highlight one of the most popular methods in video prediction in pixel space, motion-based synthesis.
@@ 3.1.4 Motion-based synthesis
There are still ample opportunities for exploration in the representation of motion. For example, GaussianPrediction Zhao et al. (2024) explores the use of 3D Gaussian representations for modeling the appearance and geometry of dynamic scenes, enabling the rendering of images for future scenarios. MoSca (Lei et al., 2024) proposes the Motion Scaffold representation, which lifts the predictions of 2D foundational models (such as depth estimation, pixel trajectories, and optical flow) to 3D and optimizes them using physics-inspired constraints. Ultimately, MoSca reconstructs the geometry and appearance of the scene with a set of dynamic Gaussians (Wu et al., 2024b). Building on video diffusion models, V3D (Chen et al., 2024g) leverages their learned world simulation capabilities to perceive the 3D world, introduces a geometrical consistency prior, and finetunes the video diffusion model into a multi-view consistent 3D generator. Seamlessly integrating new representation methods with existing frameworks presents novel challenges.
@@ 4 Stochastic Synthesis
In this section, we explore stochastic synthesis methods, including GANs and VAEs, which were originally categorized for their emphasis on modeling randomness and uncertainty in video motion. Although these models possess generative capabilities, they were originally more closely associated with stochasticity due to their primary focus on capturing the inherent variability and unpredictability of video sequences. Current research prefers to simulate the complex dynamics of the future using end-to-end generative modeling strategies rather than by introducing stochastic distributions or probabilistic models into deterministic algorithms to obtain diverse outcomes. In the following Section 5, we will discuss generative synthesis approaches, such as diffusion models and auto-regressive models, which are explicitly designed to prioritize the generation of diverse, high-quality video content. This distinction reflects the evolving focus of video synthesis research, shifting from an emphasis on stochasticity to a broader emphasis on generative capability.
@@ 4.3 Motion-Controllable Synthesis
SEINE (Chen et al., 2024e) ICLR’24 Propose a short-to-long video diffusion model, aimed at generating coherent long videos at the "story-level".
@@
CineMaster (Wang et al., 2025a) SIGGRAPH’25 A framework for 3D-aware and controllable text-to-video generation, empowering users with precise control over the scene.
@@
LivePhoto (Chen et al., 2024d) ECCV’24 Enable users to control the temporal motion of images with text, and decode motion-related instructions into videos using a generator equipped with a motion module.
@@
LVT (Rakhimov et al., 2021) VISIGRAPP’21 Reduce the computational requirements of video generative models by modeling dynamics in the latent space.
@@ 5 Generative Synthesis
The emergence and disappearance of objects introduce unpredictability, and generative models require a profound understanding of the underlying physical principles that govern the real world. Rather than relying on simplistic linear motion predictions extrapolated from historical frames, they address the challenge using sophisticated and imaginative modeling techniques. As a result, tasks such as transforming one static image into one dynamic video—often referred to as the image animation problem—have emerged as promising candidates for applying generative video prediction techniques. In general, the two mainstream branches are auto-regressive models and diffusion models. Although current auto-regressive models are inferior to diffusion models in terms of image generation quality, their structural consistency with LLMs is highly appealing for building a unified multi-modal system.
@@ 5.1 Diffusion-Based Generation
Extensions of LDM have demonstrated strong generative capabilities in video systhesis (Voleti et al., 2022). For instance, Video LDM (Blattmann et al., 2023b) leverages pre-trained image models to generate videos, enabling multi-modal, high-resolution, and long-term video synthesis. Similarly, SEINE (Chen et al., 2024e) introduces a versatile video diffusion model capable of generating transition sequences, thereby extending short clips into longer videos. Burgert et al. (2025) presents a scalable method for finetuning video diffusion models to control motion without modifying model architectures or training pipelines. It introduces a noise warping algorithm that replaces random temporal Gaussian noise with optical flow-guided warped noise, while maintaining spatial Gaussianity. Magic141 (Yi et al., 2025) decomposes the text-to-video generation task into two simpler subtasks for diffusion step distillation: text-to-image generation and image-to-video generation. It shows that image-to-video generation converges more easily than direct text-to-video generation under the same optimization setup, and explores the trade-off between computational cost and video quality.
Text-guided generation and additional information.
@@
Recent research efforts have focused on harnessing additional modalities alongside RGB images to accomplish the task of text-guided video generation Hong et al. (2022); Singer et al. (2022); Chai et al. (2023). LFDM (Ni et al., 2023) extends latent diffusion models to synthesize optical flow sequences in the latent space based on textual guidance. Seer (Gu et al., 2024) inflates Stable Diffusion (Rombach et al., 2022) along the temporal axis, enabling the model to utilize natural language instructions and reference frames to envision multiple variations of future outcomes. Emu Video (Girdhar et al., 2024) generates an image conditioned on textual guidance and extrapolates it into a video, making it adaptable to diverse textual inputs. DynamiCrafter (Xing et al., 2024) extends text-guided image animation to open-domain image scenarios. SparseCtrl (Guo et al., 2024) supports sketch-to-video generation, depth-to-video generation, and video prediction with an expanded range of input modalities. Other methods, such as PEEKABOO (Jain et al., 2024), explore interactive synthesis, aiming to unlock unprecedented applications and creative potential. VACE (Jiang et al., 2025) proposes a unified conditioning format called the video condition unit, which encodes text instructions, reference images, and masks. It incorporates a context adapter module that injects task-specific context into the model using structured temporal and spatial representations. This design ensures coherence across frames and spatial regions during generation and supports diverse applications through flexible task composition. SkyReels-A2 (Fei et al., 2025) defines the elements-to-video task, which involves generating videos by integrating arbitrary visual elements such as characters, objects, and backgrounds through reference images under the guidance of text prompts. The model demonstrates a strong ability to preserve the distinct visual details of each individual element throughout the generation process while simultaneously producing results that are coherently blended. This capability represents a significant step forward in the field by shifting the focus of video synthesis from single-condition generation to multi-condition combinatorial control. CineMaster (Wang et al., 2025a) proposes a 3D-aware, controllable text-to-video framework where users interactively define 3D control signals such as object bounding boxes and camera trajectories to guide a text-to-video diffusion model. To address the scarcity of annotated data, it introduces an automated pipeline that extracts 3D control signals from large-scale video datasets.
@@
Some works aim to achieve a more precise interpretation of textual guidance and preserve this information across the temporal dimension. MicroCinema (Wang et al., 2024b) adopts a divide-and-conquer strategy to address challenges related to appearance and temporal coherence. It employs a two-stage generation pipeline, first creating an initial image using an existing text-to-image generator, and then introducing a dedicated text-guided video generation framework for motion modeling. LivePhoto (Chen et al., 2024d) proposes a framework that incorporates motion intensity as an auxiliary factor to enhance control over motion dynamics. It further introduces a text re-weighting mechanism to emphasize motion-related descriptions, demonstrating strong performance in text-guided video synthesis tasks. I2VGen-XL (Zhang et al., 2023) utilizes static images to provide semantic and quality-related guidance, highlighting the diversity of approaches in text-guided video synthesis. SkyReels-V2 (Chen et al., 2025) introduces a multi-modal LLM to interpret narrative semantics by decomposing script-like prompts into shot-level descriptions and utilizes a specialized video captioning module to automatically annotate training data with rich cinematographic language. This methodology enables its 14B parameter model to generate videos exceeding 30 seconds in duration while exhibiting a professional cinematic style, representing the first open-source model to achieve long-form video synthesis that closely approximates the quality and coherence of film production.
@@
While the dominance of diffusion models in generative tasks is rising, some researchers still attempt to preserve the architecture of auto-regressive models (Weng et al., 2024), or develop new diffusion architectures to integrate the strengths of both paradigms (Chen et al., 2024a). Early generative video synthesis algorithms are constrained by limited data availability and model scalability, yet they lay the groundwork for integrating LLMs and diffusion models in video generation. VDT (Lu et al., 2024) represents one of the earliest efforts to incorporate a Transformer-based backbone into diffusion models for video generation. Inspired by the success of DiT (Peebles & Xie, 2023) in image synthesis, VDT replaces the conventional U-Net with a spatio-temporal Transformer that explicitly models both spatial and temporal dependencies. W.A.L.T. (Gupta et al., 2025) provides the first successful empirical demonstration of a Transformer-based backbone for jointly training image and video latent diffusion models. Recently, GAIA-1 (Hu et al., 2023a) and Sora (Brooks et al., 2024) have also leveraged the strengths of DiT to enable more creative, generalizable, and scalable video synthesis. RIVER (Davtyan et al., 2023) leverages flow matching (Lipman et al., 2023) for efficient video prediction by conditioning on a small set of past frames in the latent space of a pre-trained VQGAN (Esser et al., 2021). Collectively, these advances underscore the potential of diffusion models to produce high-quality, controllable, and diverse video content, thereby pushing the boundaries of what is achievable in computer vision and artificial intelligence.
@@
HunyuanVideo (Kong et al., 2024) is a very large open-source video generative model (13B parameters). Relying on key strategies such as data curation and image-video co-training, the model reaches a level of generation quality comparable to or even better than that of the leading closed-source models, which effectively narrows the performance gap between open-source and closed-source solutions.  Ma et al. (2025) release an even larger video generative model, Step-Video-T2V with 30B parameters, and a new benchmark for video generation, Step-Video-T2V-Eval. The model incorporates a deep compression video Variational Autoencoder for compact and efficient video representation, and leverages a flow matching approach to train a DiT (Peebles & Xie, 2022) architecture for denoising input noises into latent frames. To enhance the visual quality and mitigate artifacts, a video-based direct preference optimization method is applied. For the task of text-driven image-to-video generation, Huang et al. (2025) release Step-Video-TI2V, a model of equal scale with 30B parameters, along with its corresponding benchmark, Step-Video-TI2V-Eval. It is noteworthy that improving video generation performance does not necessarily require increasing the number of model parameters. Seaweed-7B (Seawead et al., 2025), with only 7B parameters, demonstrates generation quality comparable to that of models with tens of billions of parameters. This provides important insights into how the performance of medium-sized DiT models can be effectively enhanced without relying solely on scaling up model size. Sand-AI (2025) has open-sourced a 24B parameter video generative model, MAGI-1, that divides video frame sequences into fixed-length video chunks and adopts a chunk-wise auto-regressive diffusion framework, facilitating text-driven image-to-video generation through chunk-wise prompting. This design illustrates a promising direction in which high-fidelity video synthesis and fine-grained instruction control can be unified.
@@
While diffusion models have made significant strides in video generation, several critical challenges remain. Ensuring temporal coherence and consistency across frames is essential for achieving realism, yet remains a major challenge (Chen et al., 2024f; Xu et al., 2024). The computational efficiency and scalability of diffusion models are hindered by their resource-intensive nature, which limits their broader adoption (Peebles & Xie, 2023). Issues of controllability and interpretability persist, as textual guidance does not always align with visual outcomes, and model behaviors often remain opaque. Data availability and diversity are critical for training robust models, but acquiring comprehensive and diverse datasets remains a major bottleneck.